Find the power series representation centered at c = -3 for f(x) 2 x+5
Answer #1

Let fear = f(-3) = 1 fw0 = f(-3) = - - - - f(1) = f(-3) = 2 (x + 5)2 f (19= -4 (31+5) 3(x+ 5)² th() = -12 -- (3+5) 24 4